Slides for RMM/British Council Talk, 8 July stork: jhd/TeX/Slides/RMM/Slides1.tex Provision and Use of Information

James Davenport

Hebron & Medlock Professor of Information Technology & University Director of Information Technology & Head of Computing

University of Bath

Treasurer, European Mathematical Trust

    J.H.Davenport@bath.ac.uk

Structure

• Internal Academic Information • External Academic Information • Provision of Academic Information • Administrative Information Internal Academic Information (1) Lecture Notes “On the Web”

• Can be restricted to the institution (Internet/Intranet distinction)

• Students do like to print them out (or load them onto laptops)

• More flexible (Course Teams)

• Easier to change/correct

(but this must be made clear to the students)

• Cheaper?

– To the University

– To the student Internal Academic Information (2) Course Descriptions “On the Web”

• Useful internally (particularly in a modular system)

• Easier to keep current (but change control is still needed)

• Vital if there’s a large exchange programme (e.g. Europe’s Erasmus scheme)

• Useful in recruiting students 50% of our overseas applicants mention this Internal Academic Information (3)

Electronic Mail

The need for an “e-mail culture”, within a Department, Faculty or Institution.

• Between staff cheaper and faster than internal post can be accessed remotely Mailing lists (e.g all postgraduate admissions tutors)

• Between staff and students used sensibly it can improve academic and pastoral relationships

• Between students especially in the context of group projects Internal Academic Information (4)

Other items

• Library Catalogue Widely accessible (not just in Library) Could be much more user-friendly

• On-line telephone directory Easier to keep current (especially for functional rôles)

• Computer-based teaching/learning

– Not a universal solution

– Expensive to develop 100+ hours for 1 hour of material

– useful in the right situation External Academic Information (1)

The World-Wide Web

An enormous and ever-growing resource

Research In some subjects, e.g. High-Energy Physics, actual research communication is 100% by electronic pre-prints. Paper journals exist “as a testament to the stupidity of tenure committees”.

Teaching There is a surprising variety of material: lecture notes, handouts, hard-to-obtain material (an English–Albanian dictionary), even virtual dissection kits

http://george.lbl.gov/ITG.hm.pg.docs/
dissect/info.html

currently available in eight languages, for frogs with others planned.

External Academic Information (2)

The World-Wide Web — Myths

“The Web is for rich countries”

• The University of Zambia teaches WWW to staff and students

found on WWW in under 5 minutes

• Computing Department programs are routinely on the Web (and are used to check what an applicant might have done)

• The less information you have, the more you need a means of accessing it.

External Academic Information (3) Electronic Mail

“The Web is slow” — true, see later “The Web is for Academics” — false “The Web is unedited” — true, but irrelevant • Vital for all kinds of collaboration — books

and papers, joint research projects, software etc.

• Facilitates collaborative writing, even within an institution

• What did we do before?

Not collaborate as well

• Is a different medium from telephone or fax • External Library Catalogues

particularly national libraries

• On-line integral tables

http://http.cs.berkeley.edu/
~fateman/htest.html

• Hubble space telescope data An individual researcher can get 20 gigabytes/year

http://archive.stsci.edu/
retrieving_hst_data.html

• Video Conferencing – Used for teaching UCL→Exeter

– Used for research Bath→Southampton

Provision of Academic Information

• Computers

• Campus Networking

• National Networking

• International Networking

Computers

• University computers

• Departmental computers

• Students’ computers How are they networked?

– Wired residences?

– Dial-in modems? All these are questions of economics

Campus Networking

Is vital for all the reasons given. Today’s correct technology (especially for new buildings) is

• Within buildings:

Category 5 UTP

• Between buildings:

“Dark fibre”

Category 5 UTP as second best

• Not Cheapernet/ thin Ethernet/ BNC

National Networking

• In many developed countries, networking is largely provided by commercial (or quasicommercial) companies.

• This is not a model for developing countries to follow directly, since Universities were heavily involved in the development of national networks and in the creation of the “network culture” (and trained staff).

• The development of national networks is a matter for concerted action by Universities.

• Plan for growth! 17% per month (558% per year) is a measured growth rate.

International Networking

This is the real reason why “the Web is slow” International links never keep up with demand and the capacity of national links to fill them.

Economic These links are more expensive (under the sea; less infrastructure to reuse)

Politics Which end pays for what share of the link? Third party traffic, e.g. Belgium→Britain→ U.S.A.

The Britain↔North America link has increased by a factor of 4500 in 20 years (52% per year).

Administrative Information (1)

• How much do you need?

• How do you collect it?

• Data ownership

• Who can access it?

Administrative Information (2) How much do you need?

Enough to do your job well and no more (Principle of Parsimony)

• The UK system requires 600 bytes per student, twice a year This lets the answer questions such as “what percentage of students who, when they went to University, were living in Bath south of the river and went to Welsh universities, studied more than 20% in Welsh. Why?

• Currently at Bath, all “non-standard” students are invoiced to a single dummy code, and sorted out by hand. This makes analysis of our 5th greatest source of student revenue impossible.

Administrative Information (3) How do you collect it?

• The closer to the source of the data, the better

• On-line entry is better than batch

• It is worth investigating Optical Character/Mark Recognition (especially Mark — $10,000 several years ago)

Administrative Information (4) Data ownership

Data should be owned as centrally as possible

• Security (backup etc.)

• Consistency

• Institutional Ownership Often the Institution is responsible for data, not an individual

– 20-year old student records

– (in the case of Europe) 7-year old financial records Administrative Information (5) Who can access it?

A hard problem but not the end of the world Who can get into a filing cabinet?

20